### v1

# SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

SpotVLM: Cloud-edge Collaborative Real-time VLM
based on Context Transfer
Chen Qian1, Xinran Yu1, Zewen Huang1, Danyang Li1, Qiang Ma1,
Fan Dang2, Xuan Ding1, Guangyong Shang3, Zheng Yang1
1 Abstract
Vision–Language Models (VLMs) are increasingly deployed in real-time applications such as autonomous driving and human–computer interaction, which demand fast and reliable responses based on accurate perception. To meet these requirements, existing systems commonly employ cloud–edge collaborative architectures, such as partitioned Large Vision-Language Models (LVLMs) or task offloading strategies between Large and Small Vision-Language Models (SVLMs). However, these methods fail to accommodate cloud latency fluctuations and overlook the full potential of delayed but accurate LVLM responses. In this work, we propose a novel cloud–edge collaborative paradigm for VLMs, termed Context Transfer, which treats the delayed outputs of LVLMs as historical context to provide real-time guidance for SVLMs inference. Based on this paradigm, we design SpotVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency. Extensive experiments on three real-time vision tasks across four datasets demonstrate the effectiveness of the proposed framework. The new paradigm lays the groundwork for more effective and latency-aware collaboration strategies in future VLM systems.
2 Introduction
Figure 1: Overview of representative deployment and inference strategies for VLMs.
VLMs have witnessed rapid advancements in recent years. These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning (Xue et al. 2025), visual question answering (Fang et al. 2025), and robotic perception (Zhou et al. 2025). Notably, LVLMs (Liu et al. 2025), empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks. However, due to their substantial computational demands, these models are often deployed on the cloud servers, which inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment (Yao et al. 2024; Rang et al. 2025), enabling low-latency, on-device inference suitable for real-time scenarios like autonomous driving (Lu et al. 2025; Sima et al. 2024; Tian et al. 2024), augmented reality (Pan and Liu 2025), and smart wearable devices (Huang et al. 2024). However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs (Zhao et al. 2025).
To enable efficient and accurate real-time visual reasoning, recent studies (Hu et al. 2024b; Li et al. 2025b; Hu et al. 2024a) have explored cloud–edge collaborative architectures. As illustrated in Fig.1, existing approaches fall into two categories: (i) Model partitioning, which splits the LVLM across the cloud and edge based on hardware constraints. For example, Distributed VLM (Li et al. 2025b) deploys the visual encoder on the edge and transmits extracted features to the cloud, where they are fused with text and processed by an LLM. (ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions. For instance, ADAS (Hu et al. 2024b) offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS).
Although such methods reduce cloud-side inference workload, they still face performance bottlenecks from service latency. In video streaming scenarios, accumulated delays often cause the outputs of LVLMs arriving several frames behind the current input (e.g., producing results for frame t–d while the system is processing frame t). Consequently, although these outputs are of high quality, they are temporally outdated and cannot be effectively utilized for real-time reasoning.
To further explore more effective collaboration paradigms between large and small VLMs, we analyze the underlying factors that contribute to their performance gap and derive the following insights:
$\bullet$ Utilization of High-Quality Textual Context. Historical context plays a crucial role in generating high-quality outputs. For example, Qwen-VL (Bai et al. 2023) adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query. This context is progressively extended until the window reaches the maximum memory or model-defined limit. Such accumulation allows the model to integrate information across turns, enhancing semantic consistency and reasoning depth. More importantly, the generated outputs are recursively fed into subsequent inputs, forming a performance-enhancing feedback loop. Although small models adopt similar input strategies, their limited generative capacity often results in lower-quality historical content, which negatively affects reasoning performance. Enhancing the quality of contextual history thus emerges as a key factor in improving the reasoning ability of small models.
$\bullet$ Grounding of Semantically Salient Image Regions. The ability to align visual content with textual descriptions is critical in vision-language reasoning. However, SVLMs often struggle in this aspect, particularly when dealing with fine-grained semantics or complex scenes. Prior studies (Zheng et al. 2025; Xu et al. 2025; Zhang et al. 2025a) have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs—thereby enhancing the focus and semantic interpretability in visual reasoning. Guiding small models to focus on task-relevant image regions enhances their ability to interpret complex visual and linguistic cues, ultimately improving their reasoning performance.
Our Work: Given the inherent limitations of cloud-based LVLMs in terms of latency and response stability, we propose a key design principle in this work: Context Transfer. Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs.
Building on this principle, we introduce SpotVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
Motivated by the strong impact of historical context quality on VLM inference, we design a Context Replacement Module (CRM). Specifically, It extracts key textual segments from high-quality LVLM responses, reorganizes them based on task context, and embeds them as contextual history to replace the original low-quality history. These informative exemplars implicitly guide the small model to focus on semantically relevant visual cues that were previously overlooked, thereby enhancing cross-modal association and reasoning performance.
To exploit the grounding capabilities of large models, we introduce the Visual Focus Module (VFM), which identifies semantically salient regions from historical images based on LVLM predictions. These regions are then used to guide the SVLM: non-essential patches are masked to reduce visual tokens and improve efficiency, while attention is focused on informative areas to enhance grounding and alignment. To adapt visual focus dynamically, the edge device performs feature matching between historical and current images. Semantically aligned regions are assigned higher weights, allowing the SVLM to inherit and shift attention toward task-relevant areas in the current frame.
By providing high-quality textual history and guiding attention to semantically relevant visual regions, these modules effectively leverage the delayed outputs of LVLMs to compensate for the limited reasoning capacity of SVLMs.
The contributions of this work are summarized as follows:
  • We propose a novel architectural design paradigm, termed Context Transfer, for real-time vision-language reasoning. Instead of relying on the final predictions from LVLMs, this paradigm treats their delayed outputs as contextual priors to guide SVLMs, enabling a new form of efficient large–small model collaboration.
  • Based on this principle, we design SpotVLM with two zero-shot collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
  • Extensive experiments on four public benchmarks validate the effectiveness of SpotVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
3 Related Work
3.1 Cloud-Edge Collaborative VLMs
Figure 2: Overview of the proposed collaboration framework, SpotVLM. The system takes timestamped video streams as input and performs two parallel operations: uploading selected keyframes to the cloud-based LVLM for processing, and conducting local inference using the SVLM. Given potential cloud latency, delayed LVLM outputs are reused as historical context. These outputs guide the SVLM through two modules, Context Replacement and Visual Focus, to improve the quality of real-time predictions.
Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving (Lu et al. 2025) are shifting from modular to end-to-end pipelines (Shao et al. 2024; Li et al. 2025a), where latency sensitivity is critical and excessive delays can lead to system failure.
To enable fast and robust responses, cloud-edge collaborative architectures for VLMs have attracted increasing attention. Existing collaboration paradigms can be categorized into two types. The first category is model partitioning (Li et al. 2025b; Zhang et al. 2025b). To improve processing efficiency for LVLMs in the cloud, Distributed VLM (Li et al. 2025b) offload the visual encoder to the edge device, which encodes input images rapidly and transmits the feature embeddings to the cloud for processing and decoding. VaVLM (Zhang et al. 2025b) further reduces transmission overhead by uploading only the region-of-interest (ROI) of images. This parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings. The second category involves dynamic task allocation to different models according to predefined offloading strategies. Some systems (Hu et al. 2024b) formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency. In addition, frameworks such as LAECIPS (Hu et al. 2024a) pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models. While these approaches leverage the complementary strengths of large and small models to reduce cloud-side computation, they overlook the continued guidance that high-quality LVLM responses can provide for subsequent inference. Recognizing this underexplored potential, we propose a real-time collaborative framework that integrates delayed LVLM outputs into SVLM context to enable continual refinement and robust inference.
3.2 Historical Context in VLMs
Context refers to the sequence of prior interactions retained to guide subsequent inferences, essential for maintaining coherence in multi-turn tasks. In VLMs, this mechanism is adapted from Large Language Models (LLMs), which use fixed-length context windows to preserve relevant token sequences and enhance reasoning consistency. Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy (Sung et al. 2024) and dialogue consistency (Wei et al. 2024). Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios (Habibpour and Afghah 2025), reducing the risk of decision paralysis and oscillatory behavior. Building on this advantage, the proposed SpotVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
4 Architecture
Real-time vision-language reasoning tasks pose dual challenges in terms of processing speed and inference accuracy. To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real time. This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving at a frame rate of 1 FPS as a representative example.
4.1 Workflow
The front-facing camera continuously captures frames and sends them to the edge device. For each timestamped frame, the edge decides whether to upload it to the cloud based on the latency of the previous cloud interaction. Each frame is also processed locally by the SVLM for real-time inference. For uploaded frames, LVLM responses are handled according to latency: results within 1 second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
As illustrated in Fig.2, the output returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>. These two outputs provide guidance for subsequent frames from textual and visual perspectives, respectively. In addition, the returned result includes the timestamp of the original frame, which allows the system to assess current cloud latency and determine the temporal relevance of the result for semantic guidance.
4.2 Context Replacement
In multi-turn vision-language reasoning tasks, VLMs typically take an image and a textual question from the user as the current input. To support continuous multimodal understanding, VLMs often retain part of the historical context during inference. This insight informs the design of two novel mechanisms for enabling textual-level collaboration across LVLMs and SVLMs.
As detailed in Sec.4.1, high‑quality outputs from the large model are continuously updated and stored in an edge‑side cache as $\mathcal{H}_{T-d}^{LVLM}$:
$\mathcal{H}_{T-d}^{LVLM}\leftarrow LVLM(I_{T-d},Q_{T-d}),$ (1)
where $I_{T-d}$ and $Q_{T-d}$ denote the image and question for Frame T-d, respectively. When processing frame T, the system compares the current timestamp with the most recent timestamp T-d associated with a returned LVLM result. If the delay is within a predefined threshold $\delta$, the SVLM replaces its original output for frame T-d with the response from LVLM, and incorporates it into the historical context to guide inference on frame T.
$if\quad\Delta t=T-(T-d)\leq\delta,\quad\mathcal{H}_{T-d}^{SVLM}\leftarrow\mathcal{H}_{T-d}^{LVLM}$ (2)
Given the temporal continuity and strong correlation between adjacent frames, providing more accurate historical information increases the likelihood of correctly activating relevant tokens and suppressing irrelevant ones, thereby improving the overall precision of outputs generated by the SVLM.
Moreover, instead of directly concatenating raw questions and answers as historical context, we design an in-context learning prompt template to compress and summarize dialogue history for the SVLM, and the result is denoted as follows:
$\tilde{\mathcal{H}}=Compress(\mathcal{H}_{\leq T-1}).$ (3)
Given the limited memory capacity of edge devices and the context window constraints of lightweight models, naively stacking multi-turn inputs may lead to token truncation. By compressing the historical content, our method not only reduces the number of irrelevant tokens but also enhances attention to the semantics of prior answers for the model.
To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in the figure.
$Input_{T}^{SVLM}=\{I_{T},Q_{T},\tilde{\mathcal{H}}\}$ (4)
This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient semantic reasoning.
Figure 3: ROI-based region tracking via cosine similarity. Given ROIs of traffic cone and pedestrian from frame T–d, cosine similarity highlights related regions in frame T to guide the attention of the small model.
4.3 Visual Focus
LVLMs, equipped with strong grounding capabilities, identifies task-relevant image regions and transmits them to the edge-side SVLMs. Based on this, SVLMs can proactively focus on critical visual regions from two perspectives.
Grounding-Guided Visual Representation.
The ROIs returned by the LVLM guide the grounding of the SVLM on historical frames. Most VLMs, including the Qwen-2.5-VL-3B (Bai et al. 2025), use a Vision Transformer (ViT) that divides a resized input image into patches and encodes them into visual tokens (e.g., 1824 tokens for a 1600 × 900 image). However, many of these tokens are irrelevant to the reasoning task. To accelerate inference and enhance focus, ROIs are mapped to corresponding patches, and only task-relevant ones are selected. This selection is formally defined as:
$\mathcal{P}_{ROI}=\left\{p_{i}\in\mathcal{P}\mid p_{i}\cap ROI\neq\emptyset\right\},$ (5)
where $\mathcal{P}$ denotes the set of all image patches, and $p_{i}$ is a patch whose spatial overlap with the ROI is non-empty. In the case of multi-object recognition, the largest ROI is selected for each target category. For the case illustrated in the figure, the number of features selected for projection is reduced to 49, thereby reducing the computational cost.
nuScenes BDD100K MOT IPN Hand ACE
Micro-F1 Macro-F1 0-1 EM Micro-F1 Macro-F1 0-1 EM Acc. Acc.
SVLM Only (Local) 58.19 37.70 16.10 62.12 31.06 6.37 36.41 23.43
LVLM Only (Cloud) 45.76 32.04 20.75 49.65 25.85 20.28 23.91 61.14
Distributed VLM (Li et al. 2025b) 45.31 31.78 20.39 49.31 25.71 20.09 23.64 60.86
LAECIPS (Hu et al. 2024a) 59.84 39.64 24.29 62.47 32.14 21.05 39.40 48.71
ADAS (Hu et al. 2024b) 60.00 41.35 18.11 63.87 33.12 10.63 39.27 55.43
SpotVLM 69.96 48.19 27.32 71.19 36.26 23.88 45.11 62.75
Table 1: Performance comparison of different architectures across four datasets (values in %). All methods adopt Qwen-VL-2.5-3B as the SVLM and Qwen-VL-2.5-72B as the LVLM.
Grounding-Tracked Visual Representation.
The ROIs provided by the LVLM support cross-frame tracking, guiding the SVLM to focus on relevant regions in subsequent frames. Due to limited grounding capability, lightweight SVLMs may struggle to localize objects even when their categories are known. However, given the temporal continuity of video, ROI features from frame T-d can be matched to visual features in frame T. Let $\mathcal{P}_{T-d}^{(c)}$ denotes the set of patch indices for category $c$ on frame T-d, and $\mathbf{v}_{i}^{(T-d)}$ is the visual feature of the $i$-th patch. Thus, feature vector of object category $c$ on frame T-d is:
$\mathbf{f}^{(c)}=\frac{1}{|\mathcal{P}^{(c)}_{T-d}|}\sum_{i\in\mathcal{P}^{(c)}_{T-d}}\mathbf{v}^{(T-d)}_{i}$. (6)
We compute the cosine similarity between ROI features from frame T-d and the visual features of frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
$\mathbf{S}^{(c)}=\left[\frac{\mathbf{f}^{(c)}\cdot\mathbf{v}^{(T)}_{j}}{\|\mathbf{f}^{(c)}\|\cdot\|\mathbf{v}^{(T)}_{j}\|}\right]_{j=1}^{N},$ (7)
where $v^{(T)}_{j}$ denotes the feature of the $j$-th patch on frame T, and $N$ is the total number of patches. As shown in Fig.3, the LVLM accurately detects and localizes small objects such as traffic cones and pedestrians in frame T-d, and their mapped similarity heatmaps on frame T effectively indicate the potential locations after viewpoint shift, providing precise visual guidance for the SVLM. This cross-frame computation yields patch-level similarity scores can be aggregated into a unified similarity vector $s$, which can be defined as:
$s=\frac{1}{C}\sum_{c=1}^{C}S^{(c)}.$ (8)
Then $s$ can be used to construct the grounding-tracked visual weight $W_{s}$, formulated as follows:
$W_{s}=\mu\cdot\frac{1}{1+e^{-s}}+b,$ (9)
where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights.
To further adapt the visual representation based on semantic associations, we introduce a gated token update mechanism that enables dynamic modulation of visual tokens. The update is formulated as:
$t\leftarrow(1-\alpha)\cdot t+\alpha\cdot W_{s}\cdot t.$ (10)
Here, $t$ denotes the original visual token, and $\alpha\in[0,1]$ is a control coefficient that balances the weights between known object categories and newly emerging targets. The dynamically predicted weight $W_{s}$ modulates the visual tokens based on feature similarity, enabling adaptive token refinement under the guidance of semantic consistency.
5 Experiments
5.1 Experiments Setting
Datasets.
We evaluate the proposed architecture on three real-time tasks.
Real-time multi-object recognition. (i) nuScenes (Caesar et al. 2019): We sample camera images at 1 FPS from this autonomous driving dataset. The processed data includes 150 scenes (20 seconds each) with 10 object categories. (ii) BDD100K (MOT) (Yu et al. 2020): We sample the driving video frame at 1 FPS. The processed data includes 200 scenes (40 seconds each) with 10 object categories.
Real-time gesture recognition. IPN Hands (Benitez-Garcia et al. 2021): A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg. 120s each) to capture dynamic motion.
Real-time video frame caption. Actions for Cooking Eggs (ACE) (Shimada et al. 2012): A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
Metrics.
We adopt commonly used evaluation metrics (Durand, Mehrasa, and Mori 2019; Nam et al. 2017; Huang et al. 2017) for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM). Specifically, Micro-F1 reflects overall performance across all instances, and Macro-F1 captures class-wise performance equally. Additionally, 0-1 EM metric equals 1 only if the predicted result exactly matches the ground truth label for the current frame.
Following 0-1 EM, we model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
Note that the evaluated tasks operate under real-time constraints (1–2 FPS). If the system fails to return a valid result within one second, the output for that frame is considered empty.
Figure 4: Qualitative Results. Illustration of input questions, ground-truth answers, and system predictions from four datasets.
Implementation Details.
In the Visual Focus Module, the parameters $\mu$, $b$ and $\alpha$ are set to 0.5, 0.75, and 0.1, respectively. To balance processing speed and model performance, we retain two previous frames as context. SpotVLM consistently treats the cloud feedback as the farthest frame in the historical context of SVLM.
In addition, to evaluate the effectiveness of different collaborative architectures, we use the official pretrained versions of both the LVLMs and SVLMs without any additional fine-tuning. During inference, the outputs and response latencies of the large model are recorded for each frame to ensure fair comparison of collaboration strategies. All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
CRM GGVR GTVR Micro-F1 Macro-F1 0-1 EM
66.87 43.07 24.39
71.46 48.64 29.98
72.15 50.85 30.93
73.02 51.56 33.17
Table 2: Ablation Study. To assess the contribution of each component, we use full-frame upload and LVLM–SVLM parallel inference as the baseline, and evaluate performance improvements as modules are added incrementally.
SVLM LVLM Micro-F1 Macro-F1 0-1 EM
Baseline SpotVLM Baseline SpotVLM Baseline SpotVLM
Qwen2.5-VL-3B Qwen2.5-VL-72B 66.87 73.02 43.07 51.56 24.39 33.17
MiniCPM-o-2.6-int4 Qwen2.5-VL-72B 67.68 73.56 43.47 49.63 31.38 35.71
Qwen2.5-VL-3B GPT-4o 66.34 74.25 47.85 53.74 22.53 30.79
Table 3: Adaptability Study. Testing the SpotVLM architecture with various SVLM–LVLM combinations on nuScenes.
5.2 Overall Performance
To evaluate the effectiveness of different architectures, we assess six methods under a Wi-Fi network environment: locally deployed Qwen2.5‑VL‑3B (Bai et al. 2025), cloud-based Qwen2.5‑VL‑72B (Bai et al. 2025), Distributed VLM, LAECIPS, ADAS, and SpotVLM. The results are presented in Tab.1.
Due to cloud latency, methods that rely entirely on large models (e.g., Qwen2.5-VL-72B, Distributed VLM) yield lower F1 scores on real-time tasks than collaborative approaches with small models. However, the 0–1 EM of Qwen2.5-VL-72B remains higher than that of the small model alone, showing its superior accuracy when responses are timely. Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM, which transmits features to the cloud, slightly less effective than running full inference on the cloud model.
Collaborative architectures like LAECIPS and SpotVLM perform consistently well across all four datasets. When cloud responses are timely, they effectively leverage large-model outputs, achieving better overall metrics than Qwen2.5-VL-3B.
ADAS adopts a latency-adaptive offloading strategy. In the first three datasets, long outputs from multi-object recognition and dual-image inputs in gesture recognition result in high cloud latency. Under these conditions, ADAS relies more on the small model, yielding higher F1 scores but lower 0–1 EM compared to LAECIPS. In the fourth dataset, shorter outputs and fewer images per inference lead to more timely cloud responses. ADAS thus relies more on the output of the LVLM. Meanwhile, LAECIPS depends more on the SVLM for simple samples in this dataset, resulting in slightly lower performance than ADAS.
To ensure a fair comparison across methods, we recorded the number of cloud service invocations for ADAS and LAECIPS, and used their average values to set the interval for SpotVLM. Benefiting from its ability to effectively leverage the high-quality outputs of the LVLM under latency constraints, SpotVLM consistently achieved the best results across all tasks and collaborative strategies.
To provide a more intuitive comparison, we select one representative example from each of the four datasets for qualitative analysis. As shown in the Fig.4, for the real-time multi-object recognition task, all three methods rely on the outputs of the SVLM, with SpotVLM achieving the best performance due to guidance from the historical results of the large model. In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the cloud-based large model fails to return valid results in real time, leading to No response for ADAS. In contrast, SpotVLM leverages the history-enhanced small model to produce accurate classification results. For the video-frame captioning task, both ADAS and SpotVLM adopt the results quickly returned by the large model.
5.3 Ablation Study
We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of Contextual Replacement Module (CRM), Grounding-Guided Visual Representation (GGVR), and Grounding-Tracked Visual Representation (GTVR) in the SpotVLM framework. To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
As shown in the Tab.2, the first row corresponds to the baseline where the large and small models operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM. This baseline represents the upper-bound performance of current cloud–edge collaborative VLM architectures. Compared to this baseline, we incrementally integrate each proposed module. Results show that CRM yields the most significant performance gain. GGVR is used to reduce the number of visual tokens, while GTVR guides the SVLM to maintain consistent detection of object categories present in the historical context. Building on CRM, the integration of these two methods further enhances overall system performance across all three metrics.
5.4 Adaptability Study
To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conducted additional experiments by replacing both the small and large models. As in the ablation study, all image frames are uploaded and processed by the large model, and the upper bound is represented by a parallel inference strategy that always adopts the output of large model when available. Specifically, we adopt MiniCPM-o (Yao et al. 2024) as the SVLM on the edge side and GPT-4o (Hurst et al. 2024) as the LVLM on the cloud side. As shown in the Tab.3, the proposed SpotVLM coordination strategy consistently improves real-time inference performance across model configurations. With the continuous advancement of small and large VLMs, more powerful models can be seamlessly integrated into the SpotVLM framework to further enhance system capabilities.
Figure 5: Latency simulation. Manual simulation of total processing latency on the cloud LVLM. Red pentagrams denote the performance bounds of LVLM and SVLM; the dashed line shows the SVLM-only performance.
5.5 Latency Analysis
To evaluate the impact of cloud latency on SpotVLM, we conduct experiments on the nuScenes dataset using Qwen-2.5-VL-3B as the edge-side SVLM and preserving the outputs of Qwen-2.5-VL-72B as the cloud-side LVLM. We manually simulate various latency conditions by setting different return delays for the LVLM responses.
When the latency is 1 second, which means that the LVLM output always arrives within the real-time threshold, the system achieves performance equivalent to directly using the LVLM (denoted by the star marker). Starting from a 2-second delay, the LVLM result can only be incorporated as historical context for guiding the SVLM. Even under this condition, SpotVLM significantly outperforms the standalone SVLM (indicated by the dashed line). As latency increases, the temporal correlation between frames weakens, leading to a gradual decline in overall performance. However, SpotVLM continues to maintain a notable advantage even under a 9-second delay.
This experiment reveals that the performance upper bound is achieved when the LVLM returns results in real time for every frame, while the lower bound corresponds to using only the SVLM. The proposed collaborative architecture offers strong robustness under fluctuating latency conditions. Moreover, the results also suggest that cloud-side LVLM invocation can be dynamically scheduled. By adjusting the upload interval according to task requirements and performance targets, the system can reduce the frequency of cloud model calls without significantly compromising overall effectiveness.
6 Conclusion
In this work, we propose a novel context transfer paradigm that enables LVLMs to provide real-time guidance to SVLMs through semantic and visual contexts. Building on this paradigm, we introduce SpotVLM, a collaborative framework that integrates two corresponding modules: Context Replacement Module and Visual Focus Module. Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that SpotVLM outperforms existing collaboration methods. Moving forward, we envision integrating stronger foundation models and adaptive scheduling strategies into the collaborative framework to further optimize performance in real-world deployments.